Electrons with a kinetic energy of 6.023 x 10^4 J/mol are evolved from the surface of a metal, when exposed to a radiation of wavelength of 600 nm (photoelectric effect). The minimum amount of energy required to remove an electron from the metal atom is___. (h = 4.14*10^-15 eV-s, c = 3*10^8 m/s). 


1
Expert's answer
2021-03-18T02:51:36-0400

Absorbed energy = threshold energy + Kinetic energy of the photoelectron

hc/λ=E0

H is the Plank's constant = 6.62× 10^-34

C is the speed of light = 3×10^8

= 6.62×10^-32 ×3×10^8/600×10^-9=E0+

(6.023×10^4/6.023×10^23) J/atom

= 3.31×10^-19 = E0 +1 ×10^-19

E0 = 2.31×10^-19J

= 2.31×10^-19